OpenStudy (anonymous):

An exam consits of three true/false questions followed by five multiple choice questions each with 3 answers. How many ways can a student answer the exam if they can leave any of the questions blanks?

OpenStudy (anonymous):

They have 3 choices: true, false, and blank. And they have 3 questions where order matters. so it's just \[ 3^3 = 27\]

OpenStudy (anonymous):

Hmm I think there is a lot more ways then just 27 if that is what you're saying.

OpenStudy (anonymous):

oops you're right, I just solved for the multiple choice section give me a second

OpenStudy (anonymous):

So for multiple choice, you have 4 answers, 3 normal answers and 1 for blank. So it's: \[4^5=1024\] while for the first part of the test its: \[3^3=27 \] We multiply them, because when you must do two tasks in sequence you multiply (When you do two tasks in parallel you add) from before and we get \[ 27648 \]

OpenStudy (anonymous):

sorry for my misunderstanding

OpenStudy (anonymous):

Oooooh! No, I totally get it. I never considered blank being an option. Wow that is so much easier then trying to figure out which spots are blank at which outcomes. Haha thank you